OpenStudy (anonymous):

f(x)=x+e^(x)-1: By inspection f-inverse(0) and f-inverse(e)

OpenStudy (anonymous):

0=x+e^(x)-1

OpenStudy (anonymous):

All right, that's a little nicer: \[ 0=x+e^x-1\implies\\ 0=0+e^0-1\implies\\x=0 \]And: \[ e=x+e^x-1\implies\\ e=1+e^1-1=0+e^1\implies\\ x=1 \]Et voilá.

OpenStudy (anonymous):

how did you figure it out?

OpenStudy (anonymous):

There's no direct method. "By Inspection" (in solving, at least) generally means, 'try out what looks good' or 'look at something that seems like it could work.'

OpenStudy (anonymous):

oh yea i could do that i just could do it algebraically

OpenStudy (anonymous):

*couldn't

OpenStudy (anonymous):

Yeah, there's no nice method, at all, for these kind of things.

OpenStudy (anonymous):

Grazie!

OpenStudy (anonymous):

Di niente!

OpenStudy (anonymous):

Ciao